In a rectangle the lenght is 2cm longer than it's width & a diagonal is 10cm. What is the area??
PLease tell me the answer and how did u get it. ^^ tnx

Answer by mananth(16946)

In a rectangle the lenght is 2cm longer than it's width & a diagonal is 10cm. What is the area??
let the width be x cm
length = x+2 cm.
..
diagonal = 10 cm.
the diagonal and the sides of the rectangle form a right triangle with diagonal as the hypotenuse.
use pythagoras theorem
.
leg1^2+leg2^2= hypo^2
x^2+(x+2)^2= 10^2
x^2+x^2+4x+4 = 100
2x^2+4x+4-100=0
2x^2+4x-96=0
/2
x^2+2x-48=0
x^2+8x-6x-48=0
x(x+8)-6(x+8)=0
(x+8)(x-6)=0
x=6 which is positve
width 6 cm
length = 8 cm.
Area = L*W
Area = 6*8=48 cm^2
...
m.ananth@hotmail.ca